answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What are the symptoms of toxic shock syndrome?

TSS isn't as rare as the companies lead you to believe. It's back with vengeance and killing young women all over but it's being hidden. My 20 year old daughter died in June using Regular absorbency tampons, changing like clockwork never leaving them in over 6 hours and alternating with pads.

A friend of hers from the same school, in her class! had TSS just two months after she did!! She was 21. She read our warnings about flu symptoms and recognized it so early she removed the tampon and had her mom take her to the hospital and insisted to the Dr that it could be TSS. It saved her life.

The truth is, the deadly toxins begin to multiply within 2 hours after inserting your first tampon. Regular, or Super! When you change a tampon after even just a few hours that toxin remains inside just waiting for you to put in a new tampon. It takes up right where it left off when it comes into contact with the viscose rayon in the tampon. The longer you continue to use tampons even while changing the more concentrated the toxin becomes, Once this toxin gets into your blood stream, (it can enter with only a couple of scuffed off cells from changing a tampon) it's 10,000 times more deadly than sepsis alone.

The reason you need to use a pad at night for at least 8 hours is, you remove that rayon filled tampon and the toxin takes that 8 hours to dissipate, (die off) then in the morning you can start the deadly process all over again. Providing your tampon isn't leaving fibers behind.....

A friend of ours just had TSS, was in a coma in ICU for 3 weeks and in the hospital another 2 weeks. She wasn't using a tampon, they only found a rayon fiber from the tampon she used the prior week. She just had surgery to remove the gangrene from her toes. She is 16.

If you are young (under 28) you most likely don't yet have the antibodies for the toxin of the staphylococcus aureus bacteria, you can easily contract TSS. The symptoms are only flu symptoms! This is all my daughter had! 103.2 temp, ibuprofen brought that down, vomiting and diarrhea. That's IT! She walked into a hospital Friday and died Sunday!

If you do choose to use tampons, 100% cotton tampons are the only way to be safe. There has never been a case of TSS from the use of a 100% cotton tampon. These are not sold at Walmart, Target, Walgreen or in CVS stores. You have to order them online or go to a health food store.

95% of women have the antibodies to fight TSS by the time they are 30 years old. Please don't take chances with your life. It isn't worth it!

If you have 10 NFL games with 2 outcomes either win or lose how many different combinations are there?

I assume you mean that each team plays only one game, for a total of five games. There are 32 possible outcomes.

add-on: No, I think he/she means 10 nfl games not teams. That would be 20 teams in all. It'd probably be 2 to the 10th.. which makes it like 1024 possible outcomes.

What are the advantages and disadvantages of non probability sampling?

Advantages and disadvantages of probability sampling:

Advantages: Since it is done at random, the whole process is unbiased. This is good to use in smaller populations, of course it doesn't 100% protect from bias (depending on the question). But this option is a quicker way of achieving information.

Disadvantages: As i said before, depending on your question it cannot prevent from bias. Also, if you are sampling information from a large population this will take too much time and patience. There is also no guarantee that what these people say, represents what another group of people believe.

Non-probability sampling:

Advantages: This is more accurate because you are targeting a specific group, therefore your answers will be similar to what the rest of the population (of this group) will answer.

Disadvantages: This is more biased, because the individuals chosen are not at random. They also might not represent what another population thinks

What is the difference between the terms 'probability' and 'relative frequency'?

If an action is repeated n times and a certain event occurred b times then the ratio b/n is called the relative frequency.

Where as theoretical probability is used to determine the number of ways that the event can occur if an experiment is repeated a large number of times.

What is the probability of drawing a king of hearts from a regular deck of cards?

The probability of drawing a king of hearts from a regular deck of cards is 1 in 52 because there is only one king of hearts in the standard 52 card deck.

Mean of 85 standard deviation of 3what percent would you expect to score between 88 and 91?

mrs.sung gave a test in her trigonometry class. the scores were normally distributed with a mean of 85 and a standard deviation of 3. what percent would you expect to score between 88 and 91?

The value of a random variable are uniformly distributed between 20 and 100 the mean of this distribution is?

The value of a random variable that is uniformly distributed between 20 and 100 can be calculated by calculating the sum of numbers from 20 to 100 and dividing it by the difference between 100 and 20. The resulting mean is 58.5.

Probability of event occurring to it not occurring?

p/(1-p)

The sum of the probability of an event occurring plus it not occurring equals 1.

What are your chances of drawing a card with a prime number in FULL Deck of cards without a joker?

Assuming that the face cards are given numbers (J = 11, Q = 12, K = 13 and A = 1), then the prime numbers that are possible are 2, 3, 5, 7, 11, 13. There are six primes out of 13 possible numbers, giving a probability of 6/13, or about .462 (46.2%).

Note: This ignores the fact that there are four of each number in the deck. It's easy to show that the fours cancel out of the fraction. 24/52 = (6*4)/(13*4) = 3/13.

How was Caloris Basin created?

Probably by a big explosion caused by a meteor or a comet

A package of baseball cards contains 20 cards How many packages must be purchased in order to ensure that two cards in these packages are identical if there are a total of 550 different cards?

In the worst case, each package of cards contains 20 different cards, and each package is different than any other package. This means that, if there are 550 different cards, it would take 27.5 packages to contain all 550 cards. Since you cannot purchase a half package, you would have to purchase 28 packages to ensure that you get more than one of the same card.

How many 4s are in a deck of cards?

4 as theres onli 4 differenet suits

Spades

Diamonds

Hearts

Clubs
In a standard deck of 52 cards, there are four of everything.

How many ways can you select 3 cards from a regular deck of 52 cards?

If it matters in what order you draw them, then the number of permutations of 52 things taken three at a time is 52! - (52-3)! = 52 * 51 * 50 = 132,600.

If it does not matter in what order you draw them, the usual case, then the number of combinations of 52 things taken 3 at a time is (52! - (52-3)!) / 3! = 52 * 51 * 50 / 3 / 2 / 1 = 22,100.

What is the probability of a quarter?

I'm assuming you are asking what is the probability (P) of flipping a quarter.

This answer really depends upon how many times up are going to flip it.

If you are flipping it once, you have a 50% chance that it will land on heads and a 50% chance that it will land on tails. Either way the sum of your probabilities will add up to 1, meaning that there is a 100% chance that something will occur (see probability rules).

EX: Let H= heads and let T=tails

∑P= P(H)+P(T)=0.5+0.5=1

However, let's say you were going to flip a coin 3 times and were wanting to know what the probability of getting at least 1 tail was. You would approach the problem this way:

P( at least 1 tail)=?

Next, you want to find the compliment (the opposite of what you are starting with). So the opposite of getting one tail is getting no tails. This is the same as getting all heads.

P(no tails)=P(all heads)

P( all heads)= P(H)3 Heads is cubed because you are flipping the coin 3

= P(0.5)3 times and want all the outcomes to be heads.

= 1/8

By knowing that the outcome plus its compliment add up to equal 1 you get:

P( all heads) + P( at least 1 tail)=1

P( at least 1 tail) = 1- P( all heads)

P( at least 1 tail) = 1- 1/8

P( at least 1 tail) = 7/8

So the probability of flipping a coin 3 times and getting a least 1 tail is 7/8. In other words, it's very likely that it will land on tails one of those three times.

What cards are in a Des koala deck list?

Koala Monsters:

  • Big Koala
  • Bloodsucking Koala
  • Des Koala
  • Koala-Otter Koala
  • Sea Koala

Koala Support:

  • Des Kangaroo
  • Master of Oz
  • Tree Otter

For more information concerning the "Koala" Archetype, click on the "Related Link" below.